# Is Deep Learning finally better than Decision Trees on Tabular Data?

### v2

@@ Abstract
Tabular data is a ubiquitous data modality due to its versatility and ease of use in many real-world applications. The predominant heuristics for handling classification tasks on tabular data rely on classical machine learning techniques, as the superiority of deep learning models has not yet been demonstrated. This raises the question of whether new deep learning paradigms can surpass classical approaches. Recent studies on tabular data offer a unique perspective on the limitations of neural networks in this domain and highlight the superiority of gradient boosted decision trees (GBDTs) in terms of scalability and robustness across various datasets. However, novel foundation models have not been thoroughly assessed regarding quality or fairly compared to existing methods for tabular classification. Our study categorizes ten state-of-the-art neural models based on their underlying learning paradigm, demonstrating specifically that meta-learned foundation models outperform GBDTs in small data regimes. Although dataset-specific neural networks generally outperform LLM-based tabular classifiers, they are surpassed by an AutoML library which exhibits the best performance but at the cost of higher computational demands.
@@ 1 Introduction
Tabular data has long been one of the most common and widely used data formats, with applications spanning various fields such as healthcare (Johnson et al., 2016; Ulmer et al., 2020), finance (A. & E., 2022), and manufacturing (Chen et al., 2023), among others. Despite being a ubiquitous data modality, tabular data has only been marginally impacted by the deep learning revolution (Van Breugel & Van Der Schaar, 2024). A significant portion of the research community in tabular data mining continues to advocate for traditional machine learning methods, such as gradient-boosting decision trees (GBDTs) (Friedman, 2001; Chen & Guestrin, 2016; Prokhorenkova et al., 2018). Recent empirical studies agree that GBDTs are still competitive for tabular data (McElfresh et al., 2023). Nevertheless, an increasing segment of the community highlights the benefits of deep learning methods (Kadra et al., 2021; Gorishniy et al., 2021; Arik & Pfister, 2021; Somepalli et al., 2021; Kadra et al., 2024; Holzmüller et al., 2024).
The community remains divided on whether Deep Learning approaches are the undisputed state-of-the-art methods for tabular data (Shwartz-Ziv & Armon, 2022). To resolve this debate and determine the most effective methods for tabular data, multiple recent studies have focused on empirically comparing GBDT with Deep Learning methods (Grinsztajn et al., 2022; Borisov et al., 2022; McElfresh et al., 2023). These studies suggest that tree-based models outperform deep learning models on tabular data even after tuning neural networks.
@@
In this empirical survey paper, we address a simple question: "Is Deep Learning now state-of-the-art on tabular data, compared to GBDTs?" Providing an unbiased and empirically justified answer to this question has a significant impact on the large community of practitioners. Therefore, we designed a large-scale experimental protocol using 68 diverse OpenML datasets and 11 recent baselines, including foundation models for tabular data. In our protocol, we use 10-fold cross-validation experiments for all the datasets and fairly tune the hyperparameters of all the baselines with an equally large HPO budget.
@@
Moreover, to fully unlock a model’s potential, after HPO we refit all models on the joined training and validation set. Hence, our study provides a fair investigation of post-hyperparameter optimization. We argue that this is a crucial oversight because training on the combined dataset can provide additional information to a model, potentially improving generalization, especially, in small data regimes. Hence, in our large-scale study, the aim is to compare classical gradient-boosted decision trees to most modern deep learning model families. We classify models according to their underlying paradigm and provide a taxonomy tree in Figure 1 including tree-based models stemming from classical machine learning, in-context learning and fine-tuned models as sub-paradigms of foundation models, and dataset-specific architectures for tabular data, which we subclassify into feedforward approaches and transformer-based models. Additionally, we include the performance of an AutoML library in our study.
In summary, we employ two well-established models from traditional machine learning, namely XGBoost (Chen & Guestrin, 2016) and CatBoost (Prokhorenkova et al., 2018), and compare them with 8 deep learning architectures (Somepalli et al., 2021; Gorishniy et al., 2021; Arik & Pfister, 2021; Hollmann et al., 2023; Kim et al., 2024; Yan et al., 2024) categorized in model families w.r.t. their learning paradigm. For a fair comparison, we evaluate all models on 68 benchmark datasets from the OpenML datasets (Bischl et al., 2021). The key findings of this study are:
  • Meta-learned foundation models (Hollmann et al., 2023) outperform GBDTs in small datasets.
  • In mid-size and large datasets, XGBoost and CatBoost (Chen & Guestrin, 2016; Prokhorenkova et al., 2018) continue to be competitive against Deep Learning methods in terms of performance per training cost.
  • AutoML libraries, particularly AutoGluon (Erickson et al., 2020), deliver the highest overall performance compared to other methods but introduce a significant computational overhead.
  • Dataset-specific neural networks (e.g., FT-Transformer, SAINT, ResNet) generally outperform LLM-based tabular classifiers.
@@ 2 Related Work
Given the prevalence of tabular data in numerous areas, including healthcare, finance, psychology, and anomaly detection, as highlighted in various studies (Chandola et al., 2009; Johnson et al., 2016; Guo et al., 2017; Ulmer et al., 2020; Urban & Gates, 2021; A. & E., 2022; Van Breugel & Van Der Schaar, 2024), there has been significant research dedicated to developing algorithms that effectively address the challenges inherent in this domain. We summarize all algorithms evaluated in our study in a taxonomy tree shown in Figure 1.
@@
Classical Machine Learning.  Gradient Boosted Decision Trees (GBDTs) (Friedman, 2001), including popular implementations like XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018), are widely favored by practitioners for their robust performance on tabular datasets, and their short training times.
Table 1: Comparison with prior empirical survey works. In our study, we include 6 model families: Gradient Boosted Decision Trees (GBDT), AutoML, In-Context Learning (ICL), Fine-tuning (FT), Feed forward neural networks (FNN), and Transformer-based Models (Transformer).
@@
Protocol Model families Study Refitting Model-based HPO # Datasets GBDT AutoML ICL FT FNN Transformer # Baselines Ye et al. (2025) 300 31 Rubachev et al. (2024) 8 14 McElfresh et al. (2023) 176 19 Borisov et al. (2022) 5 20 Grinsztajn et al. (2022) 45 7 Shwartz-Ziv & Armon (2022) 11 5 Gorishniy et al. (2021) 11 11 Ours 68 13
@@
Deep Learning.  In terms of neural networks, prior work shows that meticulously searching for the optimal combination of regularization techniques in simple multilayer perceptrons (MLPs) called Regularization Cocktails (Kadra et al., 2021) can yield impressive results. Two recent papers (Kadra et al., 2021; Gorishniy et al., 2021) propose adaptations of the ResNet architecture for tabular data, demonstrating the potential of deep learning approaches in handling tabular data. This version of ResNet, originally conceived for image processing (He et al., 2016), has been effectively repurposed for tabular datasets in their research. We demonstrate that with thorough hyperparameter tuning, a ResNet model tailored for tabular data rivals the performance of transformer-based architectures. Furthermore, recent research underscores that numerical embeddings (Gorishniy et al., 2022) for tabular data are underexplored. Incorporating these embeddings into neural network architectures, including MLPs and transformer-based models, can substantially enhance performance. Additionally, novel approaches such as RealMLP (Holzmüller et al., 2024) introduce various enhancements to the standard MLP architecture. These include using robust scaling at the preprocessing stage and experimenting with alternative numerical embedding strategies. In doing so, the authors show that RealMLP surpasses other neural network models but also remains competitive with GBDT methods.
@@
Reflecting their success in various domains, transformers have also garnered attention in the tabular data domain. TabNet (Arik & Pfister, 2021), an innovative model in this area, employs attention mechanisms sequentially to prioritize the most significant features. SAINT (Somepalli et al., 2021), draws inspiration from the seminal transformer architecture (Vaswani et al., 2017). It addresses data challenges by applying attention both to rows and columns. They also offer a self-supervised pretraining phase, particularly beneficial when labels are scarce. The FT-Transformer (Gorishniy et al., 2021) stands out with its two-component structure: the Feature Tokenizer and the Transformer. The Feature Tokenizer is responsible for converting numerical and categorical features into embeddings. These embeddings are then fed into the Transformer, forming the basis for subsequent processing.
@@
Recently, a new avenue of research has emerged, focusing on the use of foundation models for tabular data. XTab (Zhu et al., 2023) utilizes shared Transformer blocks, similar to those in FT-Transformer (Gorishniy et al., 2021), followed by fine-tuning dataset-specific encoders. Another notable work, TabPFN (Hollmann et al., 2023), employs in-context learning (ICL), by leveraging sequences of labeled examples provided in the input for predictions, thereby eliminating the need for additional parameter updates after training. TP-BERTa (Yan et al., 2024), a pre-trained language model for tabular data prediction, uses relative magnitude tokenization to convert scalar numerical features into discrete tokens. TP-BERTa also employs intra-feature attention to integrate feature values with feature names. The last layer of the model is then fine-tuned on a per-dataset basis. In contrast, CARTE (Kim et al., 2024) utilizes a graph representation of tabular data and a neural network capable of capturing the context within a table. The model is then fine-tuned on a per-dataset basis.
@@
Empirical Studies.  Significant research has delved into understanding the contexts where neural networks (NNs) excel, and where they fall short (Shwartz-Ziv & Armon, 2022; Borisov et al., 2022; Grinsztajn et al., 2022). The recent study by (McElfresh et al., 2023) is highly related to ours in terms of research focus. However, the authors used only random search for tuning the hyperparameters of neural networks, whereas we employ Tree-structured Parzen Estimator (TPE) (Bergstra et al., 2011) as employed by (Gorishniy et al., 2021), which provides a more guided and efficient search strategy. Additionally, (McElfresh et al., 2023) study was limited to evaluating a maximum of 30 hyperparameter configurations, in contrast to our more extensive exploration of up to 100 configurations. Furthermore, despite using the validation set for hyperparameter optimization (HPO), they do not retrain the model on the combined training and validation data using the best-found configuration before evaluating the model on the test set. Our paper delineates from prior studies by applying a methodologically correct experimental protocol involving thorough HPO for neural networks. Moreover, Table 1 summarizes the model families evaluated in related empirical studies and highlights the differences in the evaluation protocol. To the best of our knowledge, we are the first to provide a thorough assessment of foundation models leveraging fine-tuning (FT) as a new player in the field of tabular classification (TC), and AutoML where - at the time of writing - both model families have been overlooked in most recent studies Ye et al. (2025); Rubachev et al. (2024), and compare them to other learning paradigms.
@@ 3.1 Learning with Tabular Data
A tabular dataset contains $N$ samples with $d$ features defining a $N\times d$ table. A sample $x_{i}\in\mathbb{R}^{d}$ is defined by its $d$ feature values. The features can be continuous numerical values or categorical where for the latter a common heuristic is to transform the values in numerical space. Given labels $y_{i}\in\mathcal{Y}$ being associated with the instances (rows) in the table, the task to solve is a binary or multi-class classification problem or a regression task iff $y_{i}\in\mathbb{R}$. In our study, we focus on the former. Hence, given a tabular dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, the task is to learn a prediction model $f(\cdot,\cdot)$ to minimize a classification loss function $\ell(\cdot,\cdot)$:
@@ 3.2 Experimental Setup
Datasets. In our study, we assess all the methods using OpenMLCC18 (Bischl et al., 2021), a well-established tabular benchmark in the community, which comprises 72 diverse datasets [^1] Due to memory issues encountered with several methods, we exclude four datasets from our analysis.. The datasets contain 5 to 3073 features and 500 to 100,000 instances, covering various binary and multi-class problems. The benchmark excludes artificial datasets, subsets or binarizations of larger datasets, and any dataset solvable by a single feature or a simple decision tree. For the full list of datasets used in our study, please refer to Appendix C.
Evaluation Protocol. Our evaluation employs a nested cross-validation approach. Initially, we partition the data into 10 folds. Nine of these folds are then used for hyperparameter tuning. Each hyperparameter configuration is evaluated using 9-fold cross-validation. The results from the cross-validation are used to estimate the performance of the model under a specific hyperparameter configuration.
For hyperparameter optimization, we utilize Optuna (Akiba et al., 2019), a well-known HPO library with the Tree-structured Parzen Estimator (TPE) (Bergstra et al., 2011) algorithm, the default Optuna HPO method. The optimization is constrained by a budget of either 100 trials or a maximum duration of 23 hours. Upon determining the optimal hyperparameters using Optuna, we train the model on the combined training and validation splits. We provide a detailed description of our protocol in Algorithm 1. It shows the nested-cross validation with the outer folds (lines 1-1) and inner folds (lines 1-1). In each trial (lines 1-1), the mean performance across inner folds are calculated in line 1 which is used as the objective value for Optuna in line 1. After the maximal number of trials $T$ is reached or the time budget is exceeded, we select the best hyperparameter setting in line 1. In comparison to previous studies, we now refit the model in line 1 which yields for each outer fold a performance measurement in line 1. We refer to Appendix F for an ablations study on the refitting procedure. The final performance is calculated across all outer folds in line 1. To enhance efficiency, we execute every outer fold in parallel across all datasets.
All experiments are run on NVIDIA RTX2080Ti GPUs with a memory of 11 GB. Our evaluation protocol dictates that for every algorithm, up to 68K different models will be evaluated, leading to a total of approximately 800K individual evaluations. As our study encompasses thirteen distinct methods, this methodology culminates in a substantial total of over 8M evaluations, involving more than 900K unique models.

### v3

# Tabular Data: Is Deep Learning all you need?

@@ Abstract
Tabular data represent one of the most prevalent data formats in applied machine learning, largely because they accommodate a broad spectrum of real-world problems. Existing literature has studied many of the shortcomings of neural architectures on tabular data and has repeatedly confirmed the scalability and robustness of gradient-boosted decision trees across varied datasets. However, recent deep learning models have not been subjected to a comprehensive evaluation under conditions that allow for a fair comparison with existing classical approaches. This situation motivates an investigation into whether recent deep-learning paradigms outperform classical ML methods on tabular data. Our survey fills this gap by benchmarking seventeen state-of-the-art methods, spanning neural networks, classical ML and AutoML techniques. Our empirical results over 68 diverse datasets from a well-established benchmark indicate a paradigm shift, where Deep Learning methods outperform classical approaches.
@@ 1 Introduction
Tabular data has long been one of the most common and widely used data formats, with applications spanning various fields such as healthcare (Johnson et al., 2016; Ulmer et al., 2020), finance (Nureni & Adekola, 2022), and manufacturing (Chen et al., 2023), among others. Despite being a ubiquitous data modality, tabular data has only been marginally impacted by the deep learning revolution (Van Breugel & Van Der Schaar, 2024). A significant portion of the research community in tabular data continues to advocate for traditional machine learning methods, such as gradient-boosting decision trees (GBDTs) (Friedman, 2001; Chen & Guestrin, 2016; Prokhorenkova et al., 2018; Ke et al., 2017). Recent empirical studies suggest that GBDTs are still competitive for tabular data (Shwartz-Ziv & Armon, 2022; Grinsztajn et al., 2022; McElfresh et al., 2023). Nevertheless, an increasing segment of the community highlights the benefits of deep learning methods (Kadra et al., 2021; Gorishniy et al., 2021; Arik & Pfister, 2021; Somepalli et al., 2021; Kadra et al., 2024; Holzmüller et al., 2024).
The community remains divided on whether Deep Learning approaches are the undisputed state-of-the-art methods for tabular data (Shwartz-Ziv & Armon, 2022). To resolve this debate and determine the most effective methods for tabular data, multiple recent studies have focused on empirically comparing GBDTs with Deep Learning methods (Grinsztajn et al., 2022; Borisov et al., 2022; McElfresh et al., 2023). These studies suggest that tree-based models outperform deep learning models on tabular data even after tuning neural networks.
@@
In this empirical survey paper, we address a simple question: ”Is Deep Learning now state-of-the-art on tabular data, compared to GBDTs?”. Providing an unbiased and empirically justified answer to this question has a significant impact on the large community of practitioners. Therefore, we designed a large-scale experimental protocol using 68 diverse classification OpenML datasets and 17 recent baselines, including foundation models for tabular data. We classify models according to their underlying paradigm and provide a taxonomy tree in Figure 1. In our protocol, we use 10-fold cross-validation experiments for all the datasets and fairly tune the hyperparameters of all the baselines with an equally large HPO budget. In our study, we focus on the predictive quality of models rather than interpretability. However, recent works exist that propose interpretable counterparts for the top-performing deep learning methods in our study (Kadra et al., 2024; Mueller et al., 2024).
@@
Moreover, to fully unlock a model’s potential, contrary to prior work, after HPO we refit all models on the joined training and validation set. Hence, our study provides a fair investigation of post-hyperparameter optimization. We argue that this is a crucial oversight because training on the combined dataset can further improve a model’s predictive performance and change the ranking of the models, as indicated by our empirical results. Our findings highlight a paradigm shift, where Deep Learning methods achieve state-of-the-art results and manage to outperform classical approaches.
In summary, we provide the following main insights:
  • Meta-learned foundation models and simple feed-forward neural networks outperform GBDTs in all dataset regimes.
  • Feed-forward neural networks outperform related dataset-specific architectures, and non-fine-tuned foundation models outperform fine-tuned ones.
  • Refitting on the validation dataset after performing hyperparameter optimization significantly improves predictive quality and affects the overall model rankings.
  • As an overarching contribution, to facilitate future research, we open-source our code and we release a large benchmark that includes 17 baselines run on 68 datasets, repeated 10 times with different test outer folds, for up to 100 HPO trials, yielding a total number of 8 million evaluations or a total of 11.81 GPU years.
@@ 2 Related Work
Given the prevalence of tabular data in numerous areas, including healthcare, finance, psychology, and anomaly detection, as highlighted in various studies (Chandola et al., 2009; Johnson et al., 2016; Guo et al., 2017; Ulmer et al., 2020; Urban & Gates, 2021; Nureni & Adekola, 2022; Van Breugel & Van Der Schaar, 2024), there has been significant research dedicated to developing algorithms that effectively address the challenges inherent in these domains.
@@
Classical Machine Learning.  Gradient Boosted Decision Trees (GBDTs) (Friedman, 2001), including popular implementations like XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018), are widely favored by practitioners for their robust performance on tabular datasets and their short training times.
Deep Learning.  In terms of neural networks, prior work shows that meticulously searching for the optimal combination of regularization techniques in simple multilayer perceptrons (MLPs) called Regularization Cocktails (Kadra et al., 2021) can yield impressive results. Additionally, the models in (Kadra et al., 2021; Gorishniy et al., 2021) propose adaptations of the ResNet (He et al., 2016) architecture for tabular data, demonstrating the potential of deep learning approaches in handling tabular data.
Protocol Model families
@@
Study Refitting Model-based HPO # Datasets GBDT AutoML ICL FT FNN TF # Baselines
Ye et al. (2025) × 300 × × 31
Rubachev et al. (2024) × 8 × × × 14
McElfresh et al. (2023) × × 176 × × 19
Borisov et al. (2022) × 5 × × × 20
Grinsztajn et al. (2022) × × 45 × × × 7
Shwartz-Ziv & Armon (2022) × 11 × × × × 5
Gorishniy et al. (2021) × 11 × × × 11
Ours 68 17
Table 1: Comparison with prior empirical survey works. In our study, we include 6 model families: Gradient Boosted Decision Trees (GBDT), AutoML, In-Context Learning (ICL), Fine-tuning (FT), Feed forward neural networks (FNN), and Transformer-based Models (TF).
@@
Furthermore, recent research underscores that numerical embeddings (Gorishniy et al., 2022) for tabular data are underexplored. Incorporating these embeddings into neural network architectures, including MLPs and transformer-based models, can substantially enhance performance. Moreover, novel approaches such as RealMLP (Holzmüller et al., 2024) introduce various enhancements to the standard MLP architecture. These include using robust scaling at the preprocessing stage and experimenting with alternative numerical embedding strategies. Lastly, recent research (Gorishniy et al., 2025) achieves state-of-the-art performance by combining simple feed-forward neural networks with efficient ensembling techniques effectively mimicking gradient boosted decision trees.
@@
Reflecting their success in various domains, transformers have also garnered attention in the tabular data domain. TabNet (Arik & Pfister, 2021), an innovative model in this area, employs attention mechanisms sequentially to prioritize the most significant features. SAINT (Somepalli et al., 2021) draws inspiration from the seminal transformer architecture (Vaswani et al., 2017). It addresses data challenges by applying attention both to rows and columns. They also offer a self-supervised pretraining phase, particularly beneficial when labels are scarce. The FT-Transformer (Gorishniy et al., 2021) stands out with its two-component structure: the Feature Tokenizer and the Transformer. The Feature Tokenizer is responsible for converting numerical and categorical features into embeddings. These embeddings are then fed into the Transformer, forming the basis for subsequent processing.
@@
Recently, a new avenue of research has emerged, focusing on the use of foundation models for tabular data. XTab (Zhu et al., 2023) utilizes shared Transformer blocks, similar to those in FT-Transformer (Gorishniy et al., 2021), followed by fine-tuning dataset-specific encoders. Another notable work, TabPFN (Hollmann et al., 2023), employs in-context learning (ICL), by leveraging sequences of labeled examples provided in the input for predictions, thereby eliminating the need for additional parameter updates after training. The most recent version, TabPFNv2 (Hollmann et al., 2025), addresses the limitations of the first version, handling tables with up to 10K samples, and incorporating row- and column-wise attention, improving predictive performance. TabICL (Qu et al., 2025), similar to the TabPFN models, is pretrained on millions of synthetic datasets and can scale to tables with up to 500K samples. TP-BERTa (Yan et al., 2024), a pre-trained language model for tabular data prediction, uses relative magnitude tokenization to convert scalar numerical features into discrete tokens. The last layer of the model is then fine-tuned on a per-dataset basis. In contrast, CARTE (Kim et al., 2024) utilizes a graph representation of tabular data and a neural network capable of capturing the context within a table. The model is then fine-tuned on a per-dataset basis.
@@
Empirical Studies.  Significant research has delved into understanding the contexts where neural networks (NNs) excel, and where they fall short (Shwartz-Ziv & Armon, 2022; Borisov et al., 2022; Grinsztajn et al., 2022; Rubachev et al., 2024; Ye et al., 2025). The recent study by McElfresh et al. (2023) is highly related to ours in terms of research focus. However, the authors use only random search for tuning the hyperparameters of neural networks, whereas we employ Tree-structured Parzen Estimator (TPE) (Bergstra et al., 2011) as employed by Gorishniy et al. (2021), which provides a more guided and efficient search strategy. Additionally, recent studies (McElfresh et al., 2023) are limited to evaluating a maximum of 30 hyperparameter configurations, in contrast to our more extensive exploration of up to 100 configurations. Furthermore, despite using the validation set for hyperparameter optimization (HPO), they do not retrain the model on the combined training and validation data using the best-found configuration before evaluating the model on the test set. Our paper differs from prior studies by applying a methodologically correct experimental protocol involving thorough HPO for neural networks. Lastly, Table 1 summarizes the model families evaluated in related empirical studies and highlights the differences in the evaluation protocol. To the best of our knowledge, we are the first to provide a thorough assessment of foundation models and AutoML to other learning paradigms.
@@ 3.1 Learning with Tabular Data
A tabular dataset contains $N$ samples with $d$ features defining an $N\times d$ table. A sample $x_{i}\in\mathbb{R}^{d}$ is defined by its $d$ feature values. The features can be continuous numerical values or categorical, where for the latter, a common heuristic is to transform the values into numerical space. Given labels $y_{i}\in\mathcal{Y}$ being associated with the instances (rows) in the table, the task in our study is to solve a binary or multi-class classification problem. Hence, given a tabular dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, the aim is to learn a prediction model $f(\cdot)$ to minimize a classification loss function $\ell(\cdot,\cdot)$:
@@ 3.2 Experimental Setup
Datasets. In our study, we assess all the methods using OpenMLCC18 (Bischl et al., 2021), a well-established tabular benchmark in the community, which comprises 72 diverse datasets [^1] Due to memory issues encountered with several methods, we exclude four datasets from our analysis.. The datasets contain 5 to 3073 features and 500 to 100,000 instances, covering various binary and multi-class problems. The benchmark excludes artificial datasets, subsets or binarizations of larger datasets, and any dataset solvable by a single feature or a simple decision tree. For the full list of datasets used, we kindly refer the reader to Appendix E.
Preprocessing. We use a consistent preprocessing pipeline across all methods whenever possible. By default, we apply a quantile transformation using the scikit-learn library (Pedregosa et al., 2011), and categorical features are encoded with an ordinal encoder similar to prior work (Gorishniy et al., 2021). Methods for which we do not apply this preprocessing are those that inherently require a different approach, such as TP-BERTa and CARTE, or those implemented within libraries where modifying the preprocessing pipeline is not trivial. In these cases, we use the preprocessing strategies from the original works. Regarding batch size, we do not tune it in our experiments due to memory constraints. Instead, we determine batch size heuristically, similar to the setup proposed by Chen et al. (2024), based on the number of features in the dataset. While batch sizes may vary across datasets, they remain consistent across methods.
Evaluation Protocol. Our evaluation employs a nested cross-validation approach. Initially, we partition the data into 10 folds. Nine of these folds are then used for hyperparameter tuning. Each hyperparameter configuration is evaluated using 9-fold cross-validation. The results from the cross-validation are used to estimate the performance of the model under a specific hyperparameter configuration. For hyperparameter optimization, we utilize Optuna (Akiba et al., 2019), a well-known HPO library with the Tree-structured Parzen Estimator (TPE) (Bergstra et al., 2011) algorithm, the default Optuna HPO method. The optimization is constrained by a budget of either 100 trials or a maximum duration of 23 hours, similar to prior work (Kadra et al., 2021). Upon determining the optimal hyperparameters using Optuna, we train the model on the combined training and validation splits. All experiments are run on NVIDIA RTX2080Ti GPUs with 11 GB of memory. Our evaluation protocol dictates that for every algorithm, up to 68K different models will be evaluated, leading to a total of approximately 900K individual evaluations. As our study encompasses seventeen distinct methods, this methodology culminates in a substantial total of 8M evaluations, involving 900K unique models.
A detailed description of our evaluation protocol is provided in Appendix A.1. In our study, we adhere to the official hyperparameter search spaces from the respective papers for tuning every method. As a sole exception, the early stopping procedure is performed implicitly from the HPO procedure, where the number of training iterations is a hyperparameter similar to prior work (Kadra et al., 2021). We observed that this alternative form of early stopping yields better generalization. For a detailed description of the hyperparameter search spaces of all methods included in our analysis, we refer the reader to Appendix A.